# Nemotron-H: A Family of Accurate and Efficient Hybrid Mamba-Transformer Models

### v3

@@ 1 Introduction
Attention (Vaswani et al., 2023) traditionally has been the powerhouse of large language models (LLMs). Yet, to generate one token during auto-regressive inference, the self-attention layer must perform computation that scales linearly with the number of tokens seen so far since it models interactions between all pairs of tokens in a sequence (potentially in a causal way); as a result, self-attention layers also have to store state in a KV cache during inference that is linearly proportional to the number of tokens in the sequence (Kwon et al., 2023). Recent reasoning LLMs, however, exhibit inference-time scaling, where generating more tokens at inference time can improve the quality of model responses (OpenAI, 2025; DeepSeek-AI, 2025a). Thus, attention layers can fundamentally limit overall model intelligence.
@@ 2.5 Base Model Evaluations
We also include a MATH-R1 version of the MATH task, where we provide few-shot examples generated by DeepSeek-R1 (DeepSeek-AI, 2025a). Both Nemotron-H and the Qwen 2.5 models show significant improvement on the benchmark by simply changing the prompt.
@@
We compare Nemotron-H-56B-Base directly with Qwen-2.5-72B-Base (Qwen, 2025) and Llama-3.1-70B-Base (Meta, 2024). Out of the 17 task we evaluate, Nemotron-H-56B-Base achieves the highest accuracy of the three models on 9 tasks, Qwen-2.5-72B-Base achieves the highest accuracy on 7 tasks, and Llama-3.1-70B-Base achieves the highest accuracy on 1 task; Nemotron-H-56B-Base outperforms Llama-3.1-70B-Base on all but that single task (Winogrande). We further compare Nemotron-H-56B-Base with two much larger state-of-the-art models, DeepSeek-V3-671B-Base (DeepSeek-AI, 2025b) and Llama-3.1-405B-Base (Meta, 2024); we use publicly-reported numbers for these two models on the subset of common tasks where numbers are reported. Surprisingly, Nemotron-H-56B-Base remains competitive with these models, outperforming DeepSeek-V3-671B-Base and Llama-3.1-405B-Base on 4 and 5 out of the 10 overlapping tasks respectively.
@@ 4.3 Vision-Language Benchmark Results
5 Alignment and Long Context
Finally, we study the extent to which our hybrid Mamba-Transformer models can be effectively post-trained to produce instruct and long-context variants.
We use the Nemotron-H-8B-Base model as the starting point for this study and conduct instruction-tuning and long-context extension using a pipeline consisting of three stages. In the first stage (stage1), we perform supervised fine-tuning (SFT) over a data blend consisting of 6 million code- and math-related prompts and responses, as well as other data with general instruction-following tasks. In order to improve long context performance, part of the training in this stage is done on conversations extended to 512k tokens by concatenating shorter samples. In addition, we create long-context dependencies by selecting some existing turns in the conversation and generating new turns that reference them. These referencing turns are placed randomly throughout the extended conversations. We also generate some amount of inter-related segments connected by specific topics which are placed throughout the sample. We observed that training on these longer samples (e.g. 512k tokens) improved RULER (Hsieh et al., 2024) scores on shorter sequence lengths as well (e.g. 128k).
In the second stage (stage2), we start with preference tuning on general-domain prompts instead of SFT. For preference optimization, we employ offline RPO (Sun et al., 2025). Following this step, we perform a round of RPO focused on narrow instruction following, similar to the style of prompts and responses in the IFEval Benchmark (Zhou et al., 2023). Samples in this stage are also randomly extended up to 32k tokens using the conversations from stage1, which helps retain long-context performance during this stage.
In the last stage (stage3), we use on-policy RPO, in which both preferred and dis-preffered responses are generated by the stage2 model checkpoint. For on-policy-RPO, we add additional safety data derived from AEGIS2.0 (Ghosh et al., 2025). We also increase the scale of the rewards used in this stage as we found it resulted in better scores for downstream benchmarks.
| Task |
| IFeval Prompt | 73.2 | 72.6 | 74.7 | ||||||
| IFeval Instruct | 81.2 | 80.2 | 82.5 | ||||||
| BFCL (V2) | 38.9 | 61.1 | 62.6 | ||||||
| MATH500 | 47.8 | 76.4 | 77.6 | ||||||
| MT-Bench | 7.49 | 8.04 | 7.90 | ||||||
| MMLU (0-shot generative) | 63.1 | 70.9 | 68.7 | ||||||
| MBPP | 75.1 | 79.9 | 81.0 | ||||||
| MBPP+ | 61.9 | 67.5 | 67.7 | ||||||
| HumanEval | 69.5 | 82.3 | 79.3 | ||||||
| HumanEval+ | 64.0 | 75.6 | 74.4 | ||||||
| GSM8K | 81.6 | 89.8 | 90.4 |
We conduct a thorough evaluation of the final aligned checkpoint across a diverse set of benchmarks, and compare to Llama-3.1-8B-Instruct and Qwen-2.5-7B-Instruct (Meta, 2024; Qwen, 2025). Table 9 shows Nemotron-H-8B-Instruct to be competitive with state-of-the-art instruction-tuned Transformer models. We also use the RULER benchmark (Hsieh et al., 2024) to evaluate each of the three model’s long-context capabilities (Table 10). Despite having only four self-attention layers, Nemotron-H-8B-Instruct is competitive with these models on the RULER long-context benchmark.
| Context length |
| 16,384 | 91.6 | 93.0 | 91.5 | ||||||
| 32,768 | 87.4 | 91.1 | 89.8 | ||||||
| 65,536 | 84.7 | 90.4 | 87.6 | ||||||
| 131,072 | 77.7 | 84.4 | 81.7 |
Comparison to Transformer model alignment.
The above results are consistent with our internal ablations comparing post-training of Transformer and hybrid Mamba-Transformer models pre-trained in similar ways.
Before developing Nemotron-H-8B-Instruct, we conducted smaller-scale post-training experiments on a Transformer (Nemotron-T-8B-Exp-Base) and hybrid model (Nemotron-H-8B-Exp-Base) pre-trained on identical datasets. Table 11 shows the accuracies of these two base models across selected benchmarks involving math, code, and general knowledge tasks. As for the final Nemotron-H-8B-Base and Nemotron-T-8B-Base models in Table 5, both experimental models reach comparable accuracy.
@@ 5 Alignment and Long Context
We post-trained Nemotron-T-8B-Exp-Base and Nemotron-H-8B-Exp-Base to obtain corresponding instruction-tuned models, Nemotron-T-8B-Exp-Instruct and Nemotron-H-8B-Exp-Instruct. Both models were post-trained using a multi-stage pipeline similar to the one described above for Nemotron-H-8B-Instruct. More specifically, we used:
Initial supervised fine-tuning (SFT), focusing on code and math (20 million samples), as well as general instruction-following tasks (200k examples).
Subsequent preference optimization using a dataset of 200k paired comparisons.
For preference optimization, we experimented with using multiple rounds of iterative offline RPO and multiple rounds of iterative DPO. Overall, we observed the two methods to be similar, with marginal differences between them. For the comparison in this section, we select the best final model amongst all candidate models produced through RPO or DPO; the best Nemotron-T-8B-Exp-Instruct model was produced used three rounds of iterative offline RPO and the best Nemotron-H-8B-Exp-Instruct model was produced using two rounds of iterative DPO.
We compare Nemotron-H-8B-Exp-Instruct and Nemotron-T-8B-Exp-Instruct on key alignment benchmarks in Table 12. Across these benchmarks, both models achieved comparable results. We also observed similar long-context capabilities for the two models, with both models reaching comparable RULER scores for context lengths of 131,072. These ablation experiments provide additional evidence that hybrid Mamba-Transformer models can be post-trained to yield instruction-tuned models that are on-par or better when compared to equivalent Transformer models.
@@ 6 Conclusions
Data. Abhinav Khattar, Aleksander Ficek, Amala Sanjay Deshmukh, Andrew Tao, Ayush Dattagupta, Brandon Norick*, Chengyu Dong*, Dan Su*, Daria Gitman, Evelina Bakhturina, Igor Gitman, Ilia Karmanov, Ivan Moshkov, Jane Polak Scowcroft, Jarno Seppanen, Jiaxuan You, Jocelyn Huang, John Kamalu*, Joseph Jennings*, Jupinder Parmar*, Karan Sapra, Kateryna Chumachenko, Kezhi Kong*, Lukas Voegtle, Lindsey Pavao, Markus Kliegl*, Matvei Novikov, Mehrzad Samadi, Miguel Martinez, Mostofa Patwary*, Osvald Nitski, Philipp Fischer, Pritam Gundecha, Rabeeh Karimi, Sean Narenthiran, Sanjeev Satheesh*, Seungju Han, Shrimai Prabhumoye*, Shubham Pachori, Shubham Toshniwal, Siddhartha Jain, Somshubra Majumdar, Syeda Nahida Akter*, Timo Roman, Ushnish De, Vahid Noroozi, Vitaly Kurin, Wasi Uddin Ahmad, Wei Du, Yao Xu, Yejin Choi, Ying Lin*.
@@
Alignment and long context. Adithya Renduchintala*, Ali Taghibakhshi*, Ameya Sunil Mahabaleshwarkar, Dima Rekesh*, Ellie Evans, Fei Jia, Gerald Shen*, Haifeng Qian*, Jiaqi Zeng*, Makesh Narsimhan Sreedhar, Michael Evans, Olivier Delalleau*, Prasoon Varshney, Samuel Kriman*, Shantanu Acharya*, Soumye Singhal*, Tugrul Konuk*, Yian Zhang*, Yoshi Suhara*, Zijia Chen.
@@
Software support. Adithya Renduchintala, Ali Taghibakhshi, Anna Shors, Ashwath Aithal, Balaram Buddharaju, Bobby Chen, Deepak Narayanan, Dmytro Pykhtar, Duncan Riach, Gerald Shen, Helen Ngo, Jared Casper, Jimmy Zhang, Keshav Santhanam, Lawrence McAfee, Luis Vega, Nima Tajbakhsh, Parth Chadha, Piotr Bialecki, Roger Waleffe, Sahil Jain, Terry Kong, Tyler Poon, Vijay Korthikanti, Yoshi Suhara, Zhiyu Li.
Evaluations and safety. Ameya Sunil Mahabaleshwarkar*, Christopher Parisien, David Mosallanezhad, Denys Fridman, Eileen Long, Erick Galinkin, Ewa Dobrowolska, Katherine Luna, Leon Derczynski, Marta Stepniewska-Dziubinska, Michael Evans, Roger Waleffe*, Sanjeev Satheesh*, Shaona Ghosh, Shrimai Prabhumoye, Suseella Panguluri, Syeda Nahida Akter, Varun Singh.

### v4

@@ 1 Introduction
Attention (Vaswani et al., 2023) traditionally has been the powerhouse of large language models (LLMs). Yet, to generate one token during auto-regressive inference, the self-attention layer must perform computation that scales linearly with the number of tokens seen so far since it models interactions between all pairs of tokens in a sequence (potentially in a causal way); as a result, self-attention layers also have to store state in a KV cache during inference that is linearly proportional to the number of tokens in the sequence (Kwon et al., 2023). Recent reasoning LLMs, however, exhibit inference-time scaling, where generating more tokens at inference time can improve the quality of model responses (OpenAI, 2025; DeepSeek-AI, 2025b). Thus, attention layers can fundamentally limit overall model intelligence.
@@ 2.5 Base Model Evaluations
We also include a MATH-R1 version of the MATH task, where we provide few-shot examples generated by DeepSeek-R1 (DeepSeek-AI, 2025b). Both Nemotron-H and the Qwen 2.5 models show significant improvement on the benchmark by simply changing the prompt.
@@
We compare Nemotron-H-56B-Base directly with Qwen-2.5-72B-Base (Qwen, 2025) and Llama-3.1-70B-Base (Meta, 2024). Out of the 17 task we evaluate, Nemotron-H-56B-Base achieves the highest accuracy of the three models on 9 tasks, Qwen-2.5-72B-Base achieves the highest accuracy on 7 tasks, and Llama-3.1-70B-Base achieves the highest accuracy on 1 task; Nemotron-H-56B-Base outperforms Llama-3.1-70B-Base on all but that single task (Winogrande). We further compare Nemotron-H-56B-Base with two much larger state-of-the-art models, DeepSeek-V3-671B-Base (DeepSeek-AI, 2025c) and Llama-3.1-405B-Base (Meta, 2024); we use publicly-reported numbers for these two models on the subset of common tasks where numbers are reported. Surprisingly, Nemotron-H-56B-Base remains competitive with these models, outperforming DeepSeek-V3-671B-Base and Llama-3.1-405B-Base on 4 and 5 out of the 10 overlapping tasks respectively.
@@ 4.3 Vision-Language Benchmark Results
5 Nemotron-H Reasoning Models
5.1 Nemotron-H Vs. Transformer Model Alignment
We first study the extent to which our hybrid Mamba-Transformer models can be effectively post-trained into instruction-tuned and long-context variants. We conducted experiments on Transformer (Nemotron-T-8B-Exp-Base) and hybrid (Nemotron-H-8B-Exp-Base) models pretrained on identical data resulting in two models with very similar pretraining performance as seen in Table 9.
@@ 5 Nemotron-H Reasoning Models
We then post-trained both base models into instruction-tuned variants, Nemotron-T-8B-Exp-Instruct and Nemotron-H-8B-Exp-Instruct using a multistage training procedure. In stage1, we perform supervised fine-tuning (SFT) on a blend of 6 million samples including code, math, and general instruction-following tasks. To improve long-context performance, we include training on extended conversations up to 512k tokens by concatenating shorter samples. We further enhance context by inserting references to earlier turns and introducing semantically related segments across the conversation. This improves RULER (Hsieh et al., 2024) scores even at shorter context lengths (e.g., 128k).
In stage2, we switch to preference tuning using offline RPO (Sun et al., 2025) on general-domain prompts. We then perform an additional RPO round focused on narrow instruction following, following IFEval (Zhou et al., 2023) style prompts. Samples are randomly extended to 32k tokens using conversations from stage1, preserving long-context capabilities.
In stage3, we apply on-policy RPO, generating both preferred and dispreferred completions using the stage2 checkpoint. We incorporate safety data from AEGIS2.0 (Ghosh et al., 2025) and increase reward scaling, which improves performance on downstream benchmarks.
We experimented with multiple rounds of iterative offline RPO and DPO. Both yielded similar outcomes; we report results for the best model from either method. The best Transformer model used three RPO rounds; the best hybrid model used two DPO rounds.
Table 10 shows the results. These findings support that hybrid Mamba-Transformer models can be post-trained to match or exceed Transformer performance.
@@
5.2 Training Nemotron-H Reasoning Models
Following the pilot study in which we found that mamba2-hybrid models and pure transformer models have similar post-training properties we proceed to train our Nemotron-H 8B Reasoning and Nemotron-H 47B Reasoning models.
5.2.1 Supervised Finetuning Stages
The training pipeline began with supervised fine-tuning (SFT), using curated examples that include explicit reasoning traces—enclosed in <think>...</think> tags—to guide the model through step-by-step problem-solving before reaching a final answer. These traces often represent multiple possible solution paths and encourage the model to explore alternatives and iterate, improving accuracy. However, the added verbosity also increases inference cost, especially for longer traces.
To balance this, we introduced paired examples with the reasoning stripped out, allowing the model to learn when and how to respond directly.
This dual-format training helps the model adapt fluidly to different reasoning requirements.
Stage 1: Mastering Math, Code and Science Reasoning
The first phase of fine-tuning focused on math, science, and coding—domains where explicit reasoning is especially valuable. The training data here used a 5:1 ratio of reasoning to non-reasoning samples, and some of these examples are available publicly in the Llama-Nemotron-Post-Training-Dataset.
Stage 2: Expanding Instructional Coverage, Dialogue and Safety
The second phase shifted toward instruction following, safety alignment, and multi-turn dialogue—while continuing to sample from Stage 1 to retain strong STEM performance. This dataset was more compact—about 10× smaller—and offered a balanced mix of reasoning and non-reasoning samples. This helped the model generalize across a broader range of tasks while improving control over reasoning mode switching.
5.2.2 Long-Context Training
In order to improve long-context capabilities, we do additional training on an augmented version of our Stage 2 blend, consisting of samples up to 512k tokens each. The samples in this extended blend are created by concatenating the original Stage 2 SFT conversations, as well as adding several additional types of turns, with the goal of improving the long-range memory of the model. The added turns are constructed as follows:
1.
Long-range references to existing conversations. A random short conversation that appears earlier in the concatenated sample is selected; we then generate follow-up user queries that explicitly references this turn sequence by topic (e.g., “In the part of our conversation where we discussed modern authors…”). Both the referencing questions and answers are generated with DeepSeek-R1 (DeepSeek-AI, 2025a) These reference pairs are inserted throughout conversations in both the reasoning and non-reasoning subsets of the data blend (see Figure 12). This method does not require relying on the generator model’s long context understanding, as we only need the referenced conversation as context.
2.
Multi-turn references In addition to turns referencing conversations from the short blend, we also generate multiple turns with a unified topic that may reference each other. Topics for these sequences are chosen as a combination of 3 random keywords. We then generate multiple follow-up turns that each reference one or more of these segments, requiring the model to determine which of these related segments is being referred to. All of these turns are then inserted randomly thoughout the extended conversations (see Figure 13). Similarly to the previous type, this does not require a long-context generator.
3.
Document aggregation QA. A user turn embeds a long document (up to 32k tokens) drawn from the pre-training data. The turn is followed by several questions whose answers require integrating information scattered across the long document.
4.
Common-word retrieval with distractors. To target the RULER “common-words” subtask, we insert user turns that list 30-80k tokens comprising a list of keywords interleaved with random distractors. The assistant is asked to return the most frequent words. This specialised task is used only in the non-CoT subset. It is created without the use of a language model.
We find that hybrid models can learn to operate on higher context length within just a few hundred steps of training on these 512k-token extended blends. We evaluated this capability using the RULER (Hsieh et al., 2024) benchmark in non-reasoning mode, with the model achieving an 83% RULER score on 128k sequence length after this stage.
5.2.3 Reinforcement Learning with GRPO
After SFT, we applied Group Relative Policy Optimization (GRPO) in multiple phases (Shao et al., 2024). Each phase targeted specific skills—like instruction following or tool use—by creating task-specific datasets with automatic verifiers, followed by broader fine-tuning with a general reward model.
Instruction-Following Tuning
To strengthen instruction adherence, we sampled 16,000 prompts from the LMSYS Chat dataset and paired them with IFEval-style instructions. A rule-based verifier scored outputs based on how well they satisfied each instruction, creating a reward signal that prioritized following directions with precision. We have used TRT-LLM backend for policy rollouts. IFEval RL experiments provided significant boost to IFEval capabilities while the rest of the benchmarks fluctuated slightly requiring careful checkpoint selection. We have used a learning rate of which showed consistent improvement in instruction following capabilities without regressions in other model capabilities.
General Helpfulness via Reward Model
In the final RL phase, we introduced a Qwen-32B-based reward model (scoring 92.8 on RewardBench) to improve overall response helpfulness. We trained using GRPO and prompts from HelpSteer2 Wang et al. (2024b). This final stage led to noticeable gains in output quality.
5.2.4 Controlled Reasoning Inference
Inference-time behavior can be customized using simple control tags in the system prompt:
{‘reasoning’: True}triggers reasoning mode -
{‘reasoning’: False}triggers direct-answer mode -
Omitting the tag lets the model choose
[…]
5.3 Results
5.3.1 Nemotron-H-47B-Reasoning-128K
Across benchmarks in math, coding, science, tool use, and dialogue, Nemotron-H-47B-Reasoning-128K achieves accuracy on par with or better than Llama-Nemotron Super 49B V1.0, and outperforms Qwen3 32B on all non-coding benchmarks. The model supports post-training quantization of all linear layers, enabling efficient deployment with minimal accuracy loss. We also provide the quantized checkpoint and its corresponding results to demonstrate effectiveness in practical settings.
| Benchmark |
| AIME25 | 53.0 | 54.2 | 54.2 | 55.8 | ||||||||
| MATH500 | 96.6 | 96.2 | 96.8 | 91.4 | ||||||||
| GPQA-D | 65.1 | 65.7 | 64.6 | 65.2 | ||||||||
| MBPP | 90.2 | 91.8 | 91.9 | 95.5 | ||||||||
| MBPP_PLUS | 75.1 | 79.9 | 79.9 | 81.5 | ||||||||
| LCB | 41.2 | 50.2 | 53.4 | 64.2 | ||||||||
| BFCL | 72.7 | 72.9 | 73.2 | 71.3 | ||||||||
| IFEVAL-Prompt | 81.8 | 84.5 | 84.5 | 83.9 | ||||||||
| IFEVAL-Instruction | 87.4 | 89.7 | 89.6 | 88.6 | ||||||||
| Arena Hard | 85.0 | 85.0 | 80.5 | 93.8 |
5.3.2 Nemotron-H-8B-Reasoning-128K
Checkpoint merging
We observed that the smaller 8B model struggled on internal safety benchmarks. To address this, we opted for checkpoint interpolation Wortsman et al. (2022), blending in an earlier RL checkpoint aligned from a safer long-context extended checkpoint. Checkpoint interpolation is performed by linearly interpolating model weights: . We experimented with a parameter sweep over values from 0.1 to 0.9 in increments of 0.1, and found that values around 0.5 offered a good trade-off between reasoning capabilities and safe behavior.
Across benchmarks in AIME25, science, tool use, and instruction-following, Nemotron-H-8B-Reasoning-128K achieves accuracy on par with or better than Llama-Nemotron Nano 8B V1.0. The model supports post-training quantization of all linear layers, enabling efficient deployment with minimal accuracy loss. We also provide the FP8 quantized checkpoint and its corresponding results to demonstrate effectiveness in practical settings.
| Benchmark |
| AIME25 | 51.4 | 51.7 | 46.7 | ||||||
| MATH500 | 94.8 | 94.0 | 94.0 | ||||||
| GPQA-D | 53.5 | 55.1 | 54.5 | ||||||
| MBPP | 84.1 | 86.0 | 86.2 | ||||||
| MBPP_PLUS | 71.4 | 74.1 | 73.3 | ||||||
| LCB | 52.6 | 49.5 | 44.8 | ||||||
| BFCL | 63.6 | 68.8 | 70.6 | ||||||
| IFEVAL-Prompt | 71.8 | 71.4 | 71.5 | ||||||
| IFEVAL-Instruction | 79.3 | 79.6 | 79.9 |
@@ 6 Conclusions
Nemotron-H Reasoning checkpoints described are also available in Hugging Face:
Nemotron-H-47B-Reasoning. Hugging Face.
Nemotron-H-8B-Reasoning. Hugging Face.
@@
Data. Abhinav Khattar, Aleksander Ficek, Amala Sanjay Deshmukh, Andrew Tao, Ayush Dattagupta, Brandon Norick*, Chengyu Dong*, Dan Su*, Daria Gitman, Evelina Bakhturina, Igor Gitman, Ilia Karmanov, Ivan Moshkov, Jane Polak Scowcroft, Jarno Seppanen, Jiaxuan You, Jocelyn Huang, John Kamalu*, Joseph Jennings*, Jupinder Parmar*, Karan Sapra, Kateryna Chumachenko, Kezhi Kong*, Lukas Voegtle, Lindsey Pavao, Markus Kliegl*, Matvei Novikov, Mehrzad Samadi, Miguel Martinez, Mostofa Patwary*, Osvald Nitski, Philipp Fischer, Pritam Gundecha, Rabeeh Karimi Mahabadi, Sean Narenthiran, Sanjeev Satheesh*, Seungju Han, Shrimai Prabhumoye*, Shubham Pachori, Shubham Toshniwal, Siddhartha Jain, Somshubra Majumdar, Syeda Nahida Akter*, Timo Roman, Ushnish De, Vahid Noroozi, Vitaly Kurin, Wasi Uddin Ahmad, Wei Du, Yao Xu, Yejin Choi, Ying Lin*.
@@
Alignment and long context. Adithya Renduchintala*, Ali Taghibakhshi, Ameya Sunil Mahabaleshwarkar*, Bilal Kartal*, David Mosallanezhad, Dima Rekesh*, Ellie Evans, Fei Jia*, Felipe Sores*, Gerald Shen*, Haifeng Qian*, Hoo Chang Shin, Jiaqi Zeng, Julien Veron Vialard*, Luis Vega*, Makesh Narsimhan Sreedhar, Michael Evans, Olivier Delalleau, Prasoon Varshney, Samuel Kriman*, Shantanu Acharya*, Soumye Singhal, Tugrul Konuk, Yian Zhang*, Yoshi Suhara, Zijia Chen.
@@
Software support. Adithya Renduchintala, Ali Taghibakhshi, Anna Shors, Ashwath Aithal, Balaram Buddharaju, Bobby Chen, Cyril Meurillon, David Mosallanezhad, Deepak Narayanan, Dmytro Pykhtar, Duncan Riach, Ellie Evans, Fei Jia, Felipe Sores, Gerald Shen, Helen Ngo, Jared Casper, Jimmy Zhang, Keshav Santhanam, Lawrence McAfee, Luis Vega, Michael Evans, Nima Tajbakhsh, Olivier Delalleau, Parth Chadha, Piotr Bialecki, Roger Waleffe, Sahil Jain, Terry Kong, Tyler Poon, Vijay Korthikanti, Yian Zhang, Yoshi Suhara, Zhiyu Li.
Evaluations and safety. Ameya Sunil Mahabaleshwarkar*, Christopher Parisien, David Mosallanezhad*, Denys Fridman, Eileen Long, Erick Galinkin, Ewa Dobrowolska, Katherine Luna, Leon Derczynski, Marta Stepniewska-Dziubinska, Michael Evans, Roger Waleffe*, Sanjeev Satheesh*, Shaona Ghosh, Shrimai Prabhumoye, Suseella Panguluri, Syeda Nahida Akter, Varun Singh.